A mother is pulling her son on a sled. She exerts a force of 150 N at an angle of 30°. If the total mass of the child and sled is 80 kg and the µ between the sled and the snow is -0.15, how much work is done if the sled travels 100 meters?

 


1
Expert's answer
2021-06-03T18:28:25-0400
"W_M=Fs\cos{30}=(150)(100)\cos{30}=13\ kJ\\W_f=-\mu( mg-F\sin{30})s\\=-0.15( 80(9.8)-150\sin{30})100=10.6\ kJ"